The ratio of seniors to juniors in the Math Club is 2:3. If there are 21 juniors, how many seniors are in the club?
Answer by ewatrrr(24785)


Hi
Let x represent how many seniors are in the club
Question Sets Up the Proportion
2/3 = x/21 |Cross Multiplying to solve for x
3x = 42
x = 14 seniors are in the club
CHECKING our Answer***
2/3 = 14/21